consider a polynomial ax^4-bx^2+x+5. suppose of a(-3)^4 - b(-3) + 5= 2. what is the numerical value of a(3)^4 + b(3) + 5
1
Expert's answer
2013-02-26T11:05:55-0500
Suppose, the top of the light house is H metres high. The distance to the ship:
L = H/tan(alpha)
alpha - the angle of depression
The speed of the ship equals:
V = S/t = H( 1/tan(alpha1) - 1/tan(alpha2) ) /t
In our case,
alpha1= 30, alpha2 = 45
t = 10 minutes
V = H( Sqrt[3] - 1)/ t
the ship reach the light house after time:
T = S/V = H/ (tan(alpha2) *V ) = H/V
Therefore:
T = t / ( Sqrt[3] - 1) = 10/( ( Sqrt[3] - 1) min = 13.66 minutes